OpenStudy (anonymous):

Solve for x. 2(x + 6) = -4(x - 12)

OpenStudy (eric_d):

expand first

OpenStudy (eric_d):

@Apzr

OpenStudy (anonymous):

2x+12=-4x+48 2x+4x=48-12 6x=36 x=6

OpenStudy (anonymous):

I did that but i messed up a couple steps thats why i couldnt solve it thank you